# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 6.2 Second Deterministic Equivalent
$d\beta M_{1,1}K_{AA^{T}}M_{1,1}+M_{1,3}M_{3,1}+M^{2}_{1,1}+\frac{d\alpha^{2}(1+\mathrm{tr}(K_{AA^{T}}M^{2}_{1,1}))}{1-\|\sqrt{d}\alpha K_{AA^{T}}^{\frac{1}{2}}M_{1,1}K_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2}}M_{1,1}K_{AA^{T}}M_{1,1}$
@@
$\partial_{z}M^{(\mathrm{sub})}_{1,1}=M_{1,1}^{2}+\frac{d\alpha^{2}(1+\mathrm{tr}(K_{AA^{T}}M^{2}_{1,1}))}{1-\|\sqrt{d}\alpha K_{AA^{T}}^{\frac{1}{2}}M_{1,1}K_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2}}M_{1,1}K_{AA^{T}}M_{1,1}.$
@@ 7 Concluding Remarks
In conclusion, our work extended the Dyson equation framework to handle linearizations with general correlation structures. We established a global anisotropic law for pseudo-resolvents, showing that under certain conditions, generalized trace entries of the pseudo-resolvent converge to those of the Dyson equation’s solution. Unlike much of the existing literature, our approach focuses on macroscale properties, enabling us to relax distributional assumptions and better model real-world datasets. Using the developed framework, we derive a deterministic equivalent and a Gaussian equivalence theorem for the empirical test error of random features ridge regression without relying on specific distributional assumptions about the data.
Looking forward, several open questions remain. The stability condition we use is somewhat difficult to handle, and simplifying it through direct conditions on the linearization or the underlying matrix-valued polynomial would be advantageous. Additionally, our empirical simulations (e.g., Figure 1) suggest that the result might extend beyond Lipschitz activation functions, warranting further exploration. Moreover, our assumption that the norm of the random features matrix is bounded in expectation excludes low rank spike scenario, in which the norm of the random features matrix is unbounded due to the presence of a low rank spike but the spectrum is still well-behaved. Addressing this by separating bulk and spike analysis could be an interesting direction for future research.
@@ A.1 Carathéodory-Riffen-Finsler Pseudometric
where $\mathbb{D}$ denotes the open complex unit ball of unit radius and $\mathrm{Hol}(\mathcal{D},\mathbb{D})$ denotes the set of holomorphic functions from $\mathcal{D}$ to $\mathbb{D}$ [35, 34]. The pseudometric $\alpha$ is an infinitesimal Finsler pseudometric on $\mathcal{D}$, meaning that it is non-negative, lower semicontinuous, locally bounded, and satisfies $\alpha(x,tv)=|t|\alpha(x,v)$ for every $(x,v)\in\mathcal{D}\times\mathcal{X}$ and $t\in\mathbb{R}$.
@@
The pseudometric $\alpha$ is a seminorm at each point in $\mathcal{D}$, and we interpret $\mathcal{L}(\gamma)$ as the length of the curve $\gamma$ measured with respect to $\alpha$ [35]. Then, the CRF-pseudometric $\rho$ of $\mathcal{D}$ is defined as
@@
In fact, the inequality in Proposition A.1 can be replaced by an equality when the function $f$ is a biholomorphic mapping. This means that the CRF-pseudometric is biholomorphically invariant [37]. In some sense, Proposition A.1 indicates that the CRF-pseudometric is non-expansive on the space of holomorphic functions mapping a domain onto itself.
@@
For a derivation of (26), refer to [34, Example 2].
@@ A.2 General Matrix Identities
For every $M_{1},M_{2}^{T}\in\mathbb{C}^{n\times d}$ and $z\in\mathbb{C}$ such that both $M_{1}M_{2}-zI_{n}$ and $M_{2}M_{1}^{T}-zI_{d}$ are non-singular, we have $M_{1}(M_{2}M_{1}-zI_{d})^{-1}=(M_{1}M_{2}-zI_{n})^{-1}M_{1}$.
@@ A.3 Real and Imaginary Parts of Matrices
For every $M\in\mathbb{C}^{n\times n}$, $\|\Re[M]\|\vee\|\Im[M]\|\leq\|M\|$.
@@ A.5 Block Matrices
Lemma A.10 (Block matrix inversion lemma [48, Theorem 2.1]).
@@
be a block matrix with $A\in\mathbb{C}^{n\times n}$, $B,C^{T}\in\mathbb{C}^{n\times d}$ and $D\in\mathbb{C}^{d\times d}$. If $A$ is non-singular, then $M$ is non-singular if and only if $D-CA^{-1}B$ is non-singular. In this case, the inverse of $M$ is given by
@@
Alternatively, if $D$ is non-singular, then $M$ is non-singular if and only if $A-BD^{-1}C$ is non-singular. In this case, the inverse of $M$ is given by
@@ Appendix B Proof of Technical Lemmas
Let $j,k\in\{1,2,\ldots,\ell\}$ be arbitrary. For notational convenience, let $\mathcal{C}$ be a ××$\ell\times\ell\times\gamma$-dimensional tensor and $g\sim\mathcal{N}(0,I_{\gamma})$ such that $L=\mathcal{C}(g)+\mathbb{E}L$ and $[\mathcal{C}(g)]_{j,k}=\mathcal{C}_{j,k,\alpha}g_{\alpha}$. Here, we use Einstein’s notation which means that we sum over every subscript appearing at least two times in a given expression. By Stein’s lemma,
@@
Taking the limit of the quotient of this difference with $\delta$ as it approaches 00, we get that
@@
and, consequently, $\mathbb{E}[(L-\mathbb{E}L)(L-z\Lambda-i\tau I_{\ell})^{-1}]_{j,k}=-\mathbb{E}[\mathcal{C}_{j,m,\alpha}(L-z\Lambda-i\tau I_{\ell})^{-1}_{m,a}\mathcal{C}_{a,b,\alpha}(L-z\Lambda-i\tau I_{\ell})^{-1}_{b,k}]$. Note that $\mathbb{E}[(L-\mathbb{E}L)^{T}W(L-\mathbb{E}L)]_{j,k}=\mathbb{E}[\mathcal{C}_{j,a,\alpha}g_{\alpha}W_{a,b}\mathcal{C}_{b,k,\beta}g_{\beta}]=\mathbb{E}[\mathcal{C}_{j,a,\alpha}W_{a,b}\mathcal{C}_{b,k,\alpha}]$ for every $W\in\mathbb{R}^{\ell\times\ell}$ independent of $L$. ∎
@@
Fix $z\in\mathbb{H}$, let $\{a_{j}\}_{j=1}^{d}$, $\{\hat{a}_{j}\}_{j=1}^{d}$ denote the columns of $A$ and $\hat{A}$ respectively. Suppose that $l^{T}_{j}=(a_{j}^{T},0,0,\hat{a}^{T}_{j})$ and $L_{j}=l_{j}e_{n_{\text{train}}+j}^{T}+e_{n_{\text{train}}+j}l_{j}^{T}$ for every $j\in\{1,2,\ldots,d\}$, where $\{e_{j}\}_{j=1}^{\ell}$ is the canonical basis of $\mathbb{R}^{\ell}$. In particular, we may write the linearization in (22) as $L=\mathbb{E}L+\sum_{j=1}^{d}L_{j}$. For every $j\in\{1,2,\ldots,d\}$, let $P_{j}\in\mathbb{R}^{\ell\times\ell}$ be the orthogonal matrix permuting the first and $n_{\text{train}}+j$th entries exclusively and $C_{j}\in\mathbb{R}^{(\ell-1)\times(\ell-1)}$ be the matrix cycling from position $n_{\text{train}}+j-1$ to $1$. For instance, if $v=(v_{k})_{k=1}^{\ell-1}$, then
$v^{T}C^{-1}_{j}=(v_{2},v_{3},\ldots,v_{j-1},v_{1},v_{j},v_{j+1},\ldots,v_{\ell-1}).$
We will rely heavily on a Schur complement decomposition of $(P_{j}LP_{j}-zI_{\ell})^{-1}$. For every $j\in\{1,2,\ldots,d\}$, let $l_{-j}\in\mathbb{R}^{\ell-1}$ be obtained by removing the $n_{\text{train}}+j$th entry of $l_{j}$ and $L_{-j}\in\mathbb{R}^{(\ell-1)\times(\ell-1)}$ be obtained by removing the $n_{\text{train}}+j$th columns and $n_{\text{train}}+j$th row from $L$. Define the scalar $\xi_{j}:=(1+z+l_{-j}^{T}(L_{-j}-zI_{\ell-1})^{-1}l_{-j})^{-1}$ and the matrix
$\Xi_{j}:=C_{j}(L_{-j}-zI_{\ell-1})^{-1}C_{j}-\xi_{j}C_{j}(L_{-j}-zI_{\ell-1})^{-1}l_{j}l_{j}^{T}(L_{-j}-zI_{\ell-1})^{-1}C_{j}.$
We have the following block inversion formula.
@@
$(P_{j}LP_{j}-zI_{\ell})^{-1}=\begin{bmatrix}-\xi_{j}&\xi_{j}l_{-j}^{T}(L_{-j}-zI_{\ell-1})^{-1}C_{j}\\ \xi_{j}C_{j}(L_{-j}-zI_{\ell-1})^{-1}l_{-j}&\Xi_{j}\end{bmatrix}.$
@@
$P_{j}LP_{j}=\begin{bmatrix}-1&l^{T}_{-j}C_{j}^{-1}\\ C_{j}^{-1}l_{-j}&C^{-1}_{j}L_{-j}C^{-1}_{j}\end{bmatrix}$
@@
For every $j\in\{1,2,\ldots,d\}$, let $q_{j}=l_{-j}^{T}R_{-j}l_{-j}$ and $R_{-j}:=(L_{-j}-zI_{\ell})^{-1}$. Concentration of bilinear forms is a central ingredient of many random matrix theory proof. We obtain a concentration result for $q_{j}$ by adapting [46, Lemma 4].
@@
Adapting [46, Lemma 4], there exists some absolute constants $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that
@@
$P_{j}L_{j}P_{j}=\begin{bmatrix}0&l_{-j}^{T}C^{-1}_{j}\\ C^{-1}_{j}l_{-j}&0\end{bmatrix}$
@@
$=\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}\xi_{j}l_{-j}^{T}R_{-j}l_{-j}&l_{-j}^{T}R_{-j}C_{j}-\xi_{j}l_{-j}^{T}R_{-j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\\ -\xi_{j}C_{j}^{-1}l_{-j}&\xi_{j}C^{-1}_{j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\end{bmatrix}P_{j}$
$=\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}\xi_{j}l_{-j}^{T}R_{-j}l_{-j}&-\xi_{j}l_{-j}^{T}R_{-j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\\ -\xi_{j}C^{-1}_{j}l_{-j}&\xi_{j}C^{-1}_{j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\end{bmatrix}P_{j}$
@@
$=\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}\xi_{j}\tilde{l}_{-j}^{T}R_{-j}l_{-j}&\tilde{l}_{-j}^{T}R_{-j}C_{j}-\xi_{j}\tilde{l}_{-j}^{T}R_{-j}l_{-j}l_{-j}^{T}R_{-j}C_{j}\\ -\xi_{j}C^{-1}_{j}\tilde{l}_{-j}&\xi_{j}C^{-1}_{j}\tilde{l}_{-j}l_{-j}^{T}R_{-j}C_{j}\end{bmatrix}^{2}P_{j}.$
@@
where $q_{j}=l_{-j}^{T}R_{-j}l_{-j}$, $\tilde{q}_{j}=\tilde{l}_{-j}^{T}R_{-j}\tilde{l}_{-j}$, $r_{j}=\tilde{l}_{-j}^{T}R_{-j}l_{-j}$ and
$\Psi_{j}=\begin{bmatrix}\begin{subarray}{c}q_{j}-\tilde{q}_{j}+2\xi_{j}r_{j}^{2}\end{subarray}&\begin{subarray}{c}r_{j}\tilde{l}_{-j}^{T}R_{-j}C_{j}-2\xi_{j}r_{j}^{2}l_{-j}^{T}R_{-j}C_{j}+(\tilde{q}_{j}-q_{j})l_{-j}^{T}R_{-j}C_{j}\end{subarray}\\ \begin{subarray}{c}-2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}-C^{-1}_{j}l_{-j}\end{subarray}&\begin{subarray}{c}C^{-1}_{j}(l_{-j}l_{-j}^{T}-\tilde{l}_{-j}\tilde{l}_{-j}^{T})R_{-j}C_{j}+2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}l_{-j}^{T}R_{-j}C_{j}\end{subarray}\end{bmatrix}.$
@@
$\leq\frac{1}{\Im[z]}\mathbb{E}[|q-\tilde{q}|]+2|\mathbb{E}[\xi^{2}_{1}l_{-1}^{T}R_{-1}KR_{-1}l_{-1}]|$
@@
Here, we introduced the correlation matrix $K=\mathbb{E}[l_{-1}l_{-1}^{T}]$. Using Jensen’s inequality and Cauchy Schwarz,
$\|K\|$ $\leq\|\mathbb{E}[a_{1}a_{1}^{T}]\|+\|\mathbb{E}[\hat{a}_{1}a_{1}^{T}]\|+\|\mathbb{E}[a_{1}\hat{a}_{1}^{T}]\|+\|\mathbb{E}[\hat{a}_{1}\hat{a}_{1}^{T}]\|$
$=d^{-1}(\|\mathbb{E}[AA^{T}]\|+\|\mathbb{E}[\hat{A}A^{T}]\|+\|\mathbb{E}[A\hat{A}^{T}]\|+\|\mathbb{E}[\hat{A}\hat{A}^{T}]\|)$

### v4

@@ 6.2 Second Deterministic Equivalent
$d\beta M_{1,1}K_{AA^{\top}}M_{1,1}+M_{1,3}M_{3,1}+M^{2}_{1,1}+\frac{d\alpha^{2}(1+\mathrm{tr}(K_{AA^{\top}}M^{2}_{1,1}))}{1-\|\sqrt{d}\alpha K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}K_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2}}M_{1,1}K_{AA^{\top}}M_{1,1}$
@@
$\partial_{z}M^{(\mathrm{sub})}_{1,1}=M_{1,1}^{2}+\frac{d\alpha^{2}(1+\mathrm{tr}(K_{AA^{\top}}M^{2}_{1,1}))}{1-\|\sqrt{d}\alpha K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}K_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2}}M_{1,1}K_{AA^{\top}}M_{1,1}.$
@@ 7 Concluding Remarks
In conclusion, our work extends the Dyson equation framework to handle linearizations with general correlation structures. We establish a global anisotropic law for pseudo-resolvents, showing that under suitable conditions, generalized trace entries of the pseudo-resolvent converge to those of the solution of the Dyson equation. In contrast with much of the existing literature, our approach focuses on the macroscale properties of structured linearizations, allowing us to relax the usual independence assumptions and to accommodate correlated block structures. The conditions we impose include standard boundedness assumptions, a technical stability condition on the linearization, and a Gaussian design assumption on its random components. Within this framework, we derive a deterministic equivalent and a Gaussian equivalence theorem for the empirical test error of random-features ridge regression, where the training and test data are deterministic and the random weight matrix is obtained by applying an entrywise Lipschitz function to a matrix with independent standard Gaussian entries.
Looking forward, several open questions remain. While the boundedness and stability conditions are somewhat fundamental to our analysis, the Gaussian design assumption could likely be relaxed. Doing so would require replacing the Gaussian concentration arguments used here with more general probabilistic tools to ensure that the error terms vanish asymptotically and that the generalized trace entries of the pseudo-resolvent concentrate around their expectations. The current stability condition, though essential for our proofs, is technical and somewhat difficult to handle. Simplifying it through direct conditions on the linearization or the underlying matrix-valued polynomial would be highly valuable.. Additionally, our empirical simulations (e.g., Figure 1) suggest that the result might extend beyond Lipschitz activation functions, warranting further exploration. Moreover, our assumption that the norm of the random features matrix is bounded in expectation excludes low rank spike scenario, in which the norm of the random features matrix is unbounded due to the presence of a low rank spike but the spectrum is still well-behaved. Addressing this by separating bulk and spike analysis could be an interesting direction for future research.
@@ A.1 Carathéodory-Riffen-Finsler Pseudometric
where $\mathbb{D}$ denotes the open complex unit ball of unit radius and $\mathrm{Hol}(\mathcal{D},\mathbb{D})$ denotes the set of holomorphic functions from $\mathcal{D}$ to $\mathbb{D}$ [harris_fixed_2003, HARRIS1979345]. The pseudometric $\alpha$ is an infinitesimal Finsler pseudometric on $\mathcal{D}$, meaning that it is non-negative, lower semicontinuous, locally bounded, and satisfies $\alpha(x,tv)=|t|\alpha(x,v)$ for every $(x,v)\in\mathcal{D}\times\mathcal{X}$ and $t\in\mathbb{R}$.
@@
The pseudometric $\alpha$ is a seminorm at each point in $\mathcal{D}$, and we interpret $\mathcal{L}(\gamma)$ as the length of the curve $\gamma$ measured with respect to $\alpha$ [harris_fixed_2003]. Then, the CRF-pseudometric $\rho$ of $\mathcal{D}$ is defined as
@@
In fact, the inequality in Proposition A.1 can be replaced by an equality when the function $f$ is a biholomorphic mapping. This means that the CRF-pseudometric is biholomorphically invariant [helton2007operatorvalued]. In some sense, Proposition A.1 indicates that the CRF-pseudometric is non-expansive on the space of holomorphic functions mapping a domain onto itself.
@@
For a derivation of (26), refer to [HARRIS1979345, Example 2].
@@ A.2 General Matrix Identities
For every $M_{1},M_{2}^{\top}\in\mathbb{C}^{n\times d}$ and $z\in\mathbb{C}$ such that both $M_{1}M_{2}-zI_{n}$ and $M_{2}M_{1}^{\top}-zI_{d}$ are non-singular, we have $M_{1}(M_{2}M_{1}-zI_{d})^{-1}=(M_{1}M_{2}-zI_{n})^{-1}M_{1}$.
@@ A.3 Real and Imaginary Parts of Matrices
For every $\in\mathbb{C}^{n\times n}, [M][M$$\|\Re[M$$\|\vee\|\Im$$\|\leq\|M\|$.
@@ A.5 Block Matrices
The invertibility of block matrices is closely related to the invertibility of their so-called Schur complements. If $M$ is a block matrix of the form
$M=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$
with $A\in\mathbb{C}^{n\times n}$, $B,C^{\top}\in\mathbb{C}^{n\times d}$ and $D\in\mathbb{C}^{d\times d}$, then the Schur complement of $A$ in $M$ is defined as $D-CA^{-1}B$ provided that $A$ is non-singular. Similarly, the Schur complement of $D$ in $M$ is defined as $A-BD^{-1}C$ provided that $D$ is non-singular. The following lemma, known as the block matrix inversion lemma, relates the invertibility of $M$ to that of its Schur complements.
Lemma A.9 (Block matrix inversion lemma [block_matrices, Theorem 2.1]).
@@
be a block matrix with $A\in\mathbb{C}^{n\times n}$, $B,C^{\top}\in\mathbb{C}^{n\times d}$ and $D\in\mathbb{C}^{d\times d}$. If $A$ is non-singular, then $M$ is non-singular if and only if the Schur complement of $A$, namely $D-CA^{-1}B$, is non-singular. In this case, the inverse of $M$ is given by
@@
Alternatively, if $D$ is non-singular, then $M$ is non-singular if and only if the Schur complement of $D$, namely $A-BD^{-1}C$, is non-singular. In this case, the inverse of $M$ is given by
@@
A.6 Positivity-Preserving Linear Maps on Matrices
We record some observations about positivity-preserving linear maps on matrices. Recall that a linear map $\mathcal{S}:\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}$ is positivity-preserving if $\mathcal{S}(M)\succeq 0$ whenever $M\succeq 0$.
Lemma A.10.
Suppose that $\mathcal{S}:\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}$ is a positivity-preserving linear map. Then, $\mathcal{S}$ maps self-adjoint matrices to self-adjoint matrices.
Proof.
Let $M\in\mathbb{C}^{n\times n}$ be self-adjoint. We may write $M=M_{1}-M_{2}$ where $M_{1},M_{2}\succeq 0$. Since $\mathcal{S}$ is linear and positivity-preserving, $\mathcal{S}(M)=\mathcal{S}(M_{1})-\mathcal{S}(M_{2})$ is the difference of two self-adjoint matrices, and hence self-adjoint. ∎
Lemma A.11.
Suppose that $\mathcal{S}:\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}$ is a positivity-preserving linear map. Then, for any $M\in\mathbb{C}^{n\times n}$, $(\mathcal{S}(M))^{\ast}=\mathcal{S}(M^{\ast})$. In particular, it follows that $\Im[\mathcal{S}(M)]=\mathcal{S}(\Im[M])$ and $\Re[\mathcal{S}(M)]=\mathcal{S}(\Re[M])$.
Proof.
Decompose $M=\Re[M]+i\Im[M]$. By Lemma A.10, $\mathcal{S}(\Re[M])$ and $\mathcal{S}(\Im[M])$ are self-adjoint. Hence, $(\mathcal{S}(M))^{\ast}=\mathcal{S}(\Re[M])-i\mathcal{S}(\Im[M])=\mathcal{S}(M^{\ast})$. ∎
@@ Appendix B Proof of Technical Lemmas
Let $j,k\in\{1,2,\ldots,\ell\}$ be arbitrary. For notational convenience, let $\mathcal{C}$ be a $\ell\times\ell\times\gamma$-dimensional tensor and $g\sim\mathcal{N}(0,I_{\gamma})$ such that $L=\mathcal{C}(g)+\mathbb{E}L$ and $\mathcal{C}(g)]_{j,k}=\mathcal{C}_{j,k,\alpha}g_{\alpha}$. Here, we use Einstein’s notation which means that we sum over every subscript appearing at least two times in a given expression. By Stein’s lemma,
@@
Taking the limit of the quotient of this difference with $\delta$ as it approaches 0, we get that
@@
and, consequently, $\mathbb{E}[(L-\mathbb{E}L)(L-z\Lambda-i\tau I_{\ell})^{-1}]_{j,k}=-\mathbb{E}[\mathcal{C}_{j,m,\alpha}(L-z\Lambda-i\tau I_{\ell})^{-1}_{m,a}\mathcal{C}_{a,b,\alpha}(L-z\Lambda-i\tau I_{\ell})^{-1}_{b,k}]$. Note that $\mathbb{E}[(L-\mathbb{E}L)^{\top}W(L-\mathbb{E}L)]_{j,k}=\mathbb{E}[\mathcal{C}_{j,a,\alpha}g_{\alpha}W_{a,b}\mathcal{C}_{b,k,\beta}g_{\beta}]=\mathbb{E}[\mathcal{C}_{j,a,\alpha}W_{a,b}\mathcal{C}_{b,k,\alpha}]$ for every $W\in\mathbb{R}^{\ell\times\ell}$ independent of $L$. ∎
@@
Fix $z\in\mathbb{H}$, let $\{a_{j}\}_{j=1}^{d}$, $\{\hat{a}_{j}\}_{j=1}^{d}$ denote the columns of $A$ and $\hat{A}$ respectively. Suppose that $l^{\top}_{j}=(a_{j}^{\top},0,0,\hat{a}^{\top}_{j})$ and $L_{j}=l_{j}e_{n_{\text{train}}+j}^{\top}+e_{n_{\text{train}}+j}l_{j}^{\top}$ for every $j\in\{1,2,\ldots,d\}$, where $\{e_{j}\}_{j=1}^{\ell}$ is the canonical basis of $\mathbb{R}^{\ell}$. In particular, we may write the linearization in (22) as $L=\mathbb{E}L+\sum_{j=1}^{d}L_{j}$. For every $j\in\{1,2,\ldots,d\}$, let $P_{j}\in\mathbb{R}^{\ell\times\ell}$ be the orthogonal matrix permuting the first and $n_{\text{train}}+j$th entries exclusively and $C_{j}\in\mathbb{R}^{(\ell-1)\times(\ell-1)}$ be the matrix cycling from position $n_{\text{train}}+j-1$ to $1$. For instance, if $v=(v_{k})_{k=1}^{\ell-1}$, then
$v^{\top}C^{-1}_{j}=(v_{2},v_{3},\ldots,v_{j-1},v_{1},v_{j},v_{j+1},\ldots,v_{\ell-1}).$
We will rely heavily on a Schur complement decomposition of $(P_{j}LP_{j}-zI_{\ell})^{-1}$. For every $j\in\{1,2,\ldots,d\}$, let $l_{-j}\in\mathbb{R}^{\ell-1}$ be obtained by removing the $n_{\text{train}}+j$th entry of $l_{j}$ and $L_{-j}\in\mathbb{R}^{(\ell-1)\times(\ell-1)}$ be obtained by removing the $n_{\text{train}}+j$th columns and $n_{\text{train}}+j$th row from $L$. Define the scalar $\xi_{j}:=(1+z+l_{-j}^{\top}(L_{-j}-zI_{\ell-1})^{-1}l_{-j})^{-1}$ and the matrix
$\Xi_{j}:=C_{j}(L_{-j}-zI_{\ell-1})^{-1}C_{j}-\xi_{j}C_{j}(L_{-j}-zI_{\ell-1})^{-1}l_{j}l_{j}^{\top}(L_{-j}-zI_{\ell-1})^{-1}C_{j}.$
We have the following block matrix inversion formula.
@@
$(P_{j}LP_{j}-zI_{\ell})^{-1}=\begin{bmatrix}-\xi_{j}&\xi_{j}l_{-j}^{\top}(L_{-j}-zI_{\ell-1})^{-1}C_{j}\\ \xi_{j}C_{j}(L_{-j}-zI_{\ell-1})^{-1}l_{-j}&\Xi_{j}\end{bmatrix}$.
@@
$P_{j}LP_{j}=\begin{bmatrix}-1&l^{\top}_{-j}C_{j}^{-1}\\ C_{j}^{-1}l_{-j}&C^{-1}_{j}L_{-j}C^{-1}_{j}\end{bmatrix}$
@@
For every $j\in\{1,2,\ldots,d\}$, let $q_{j}=l_{-j}^{\top}R_{-j}l_{-j}$ and $R_{-j}:=(L_{-j}-zI_{\ell})^{-1}$. Concentration of bilinear forms is a central ingredient of many random matrix theory proof. We obtain a concentration result for $q_{j}$ by adapting [louart_random-matrix-approach, Lemma 4].
@@
Adapting [louart_random-matrix-approach, Lemma 4], there exists some absolute constants $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that
@@
$P_{j}L_{j}P_{j}=\begin{bmatrix}0&l_{-j}^{\top}C^{-1}_{j}\\ C^{-1}_{j}l_{-j}&0\end{bmatrix}$
@@
$=\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}\xi_{j}l_{-j}^{\top}R_{-j}l_{-j}&l_{-j}^{\top}R_{-j}C_{j}-\xi_{j}l_{-j}^{\top}R_{-j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\\ -\xi_{j}C_{j}^{-1}l_{-j}&\xi_{j}C^{-1}_{j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{bmatrix}P_{j}$
$=\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}\xi_{j}l_{-j}^{\top}R_{-j}l_{-j}&-\xi_{j}l_{-j}^{\top}R_{-j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\\ -\xi_{j}C^{-1}_{j}l_{-j}&\xi_{j}C^{-1}_{j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{bmatrix}P_{j}$
@@
$=\sum_{j=1}^{d}P_{j}\mathbb{E}\begin{bmatrix}\xi_{j}\tilde{l}_{-j}^{\top}R_{-j}l_{-j}&\tilde{l}_{-j}^{\top}R_{-j}C_{j}-\xi_{j}\tilde{l}_{-j}^{\top}R_{-j}l_{-j}l_{-j}^{\top}R_{-j}C_{j}\\ -\xi_{j}C^{-1}_{j}\tilde{l}_{-j}&\xi_{j}C^{-1}_{j}\tilde{l}_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{bmatrix}^{2}P_{j}.$
@@
where $q_{j}=l_{-j}^{\top}R_{-j}l_{-j}$, $\tilde{q}_{j}=\tilde{l}_{-j}^{\top}R_{-j}\tilde{l}_{-j}$, $r_{j}=\tilde{l}_{-j}^{\top}R_{-j}l_{-j}$ and
$\Psi_{j}=\begin{bmatrix}\begin{subarray}{c}q_{j}-\tilde{q}_{j}+2\xi_{j}r_{j}^{2}\end{subarray}&\begin{subarray}{c}r_{j}\tilde{l}_{-j}^{\top}R_{-j}C_{j}-2\xi_{j}r_{j}^{2}l_{-j}^{\top}R_{-j}C_{j}+(\tilde{q}_{j}-q_{j})l_{-j}^{\top}R_{-j}C_{j}\end{subarray}\\ \begin{subarray}{c}-2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}-C^{-1}_{j}l_{-j}\end{subarray}&\begin{subarray}{c}C^{-1}_{j}(l_{-j}l_{-j}^{\top}-\tilde{l}_{-j}\tilde{l}_{-j}^{\top})R_{-j}C_{j}+2\xi_{j}r_{j}C^{-1}_{j}\tilde{l}_{-j}l_{-j}^{\top}R_{-j}C_{j}\end{subarray}\end{bmatrix}$.
@@
$\leq\frac{1}{\Im[z]}\mathbb{E}[|q-\tilde{q}|]+2|\mathbb{E}[\xi^{2}_{1}l_{-1}^{\top}R_{-1}KR_{-1}l_{-1}]|$
@@
Here, we introduced the correlation matrix $K=\mathbb{E}[l_{-1}l_{-1}^{\top}]$. Using Jensen’s inequality and Cauchy Schwarz,
$\|K\|$ $\leq\|\mathbb{E}[a_{1}a_{1}^{\top}]\|+\|\mathbb{E}[\hat{a}_{1}a_{1}^{\top}]\|+\|\mathbb{E}[a_{1}\hat{a}_{1}^{\top}]\|+\|\mathbb{E}[\hat{a}_{1}\hat{a}_{1}^{\top}]\|$
$=d^{-1}(\|\mathbb{E}[AA^{\top}]\|+\|\mathbb{E}[\hat{A}A^{\top}]\|+\|\mathbb{E}[A\hat{A}^{\top}]\|+\|\mathbb{E}[\hat{A}\hat{A}^{\top}]\|)$